OpenStudy (anonymous):

Solve for y, then find the value of y when given x = –2. 6x = 7 – 4y a. –12 b. 7/4 (seven over four.) c. 19/4 (nineteen over four.) d. 24

OpenStudy (anonymous):

6x-7/-4=x

OpenStudy (anonymous):

Sorry, its: 6x-7/-4=y

OpenStudy (anonymous):

Then subs x for -2

OpenStudy (anonymous):

-19 / -4 = y ?

OpenStudy (anonymous):

omg thank you i got it c:

OpenStudy (anonymous):

6(-2)-7/-4=

OpenStudy (anonymous):

19/4

OpenStudy (anonymous):

Negative/negative=positive

OpenStudy (anonymous):

Yes C